find the pH of a .155 M HClO2 solution, ka=.011 for HClO2
Answer #1

For HClO2 Ka = 0.011

Ka = [H+][A-]/[HA]

H = x = H

HA = M-x = 0.155-x

0.011 = x*x/(0.155-x)

x = H = 0.03615

pH = -log(0.03615) =

pH = 1.44189